OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (campbell_st):

use the vertex to find the equation and a =1 (x +6)^2 = 4(y + 1) then x^2 + 12x + 36 = 4y + 4

OpenStudy (anonymous):

is x^2 + 12x + 36 = 4y + 4 the answer ?